OpenStudy (anonymous):

solve for y x+2y=5 answer in this format y=(__)/__

OpenStudy (paki):

simple... leave "y" alone...

OpenStudy (anonymous):

Divide by 2 on both sides

OpenStudy (anonymous):

y=5/2-x/2?

OpenStudy (paki):

no... do like this.... like 2y=5-x or y=5-x/2

OpenStudy (anonymous):

and then you get y = 5-x/2

OpenStudy (anonymous):

@ruben12345, did that help?

OpenStudy (anonymous):

what @paki said